OpenStudy (jennychan12):

Using the substitution u = 2x+1, the integral from 0 to 2 of sqrt(2x+1)dx is ...? (set up integral; do not evaluate)

OpenStudy (jennychan12):

\[\int\limits_{0}^{2} \sqrt{2x+1}dx\] when u = 2x+1

OpenStudy (jennychan12):

do the limits change?

OpenStudy (jennychan12):

cuz i got \[\frac{ 1 }{ 2 } \int\limits_{0}^{2}\sqrt{u}du \] as the answer

OpenStudy (abb0t):

That's fine. Go on.

OpenStudy (jennychan12):

ok that's all i needed.